Claim amounts modelled as an exponential random variable with the mean k1000.




(a) Calculate the probability that a random variable selected claim amount is greater than k5000.




(b) Calculate the probability that a randomly selected claim amount is greater than k5000 given that it is greater than k1000

1
Expert's answer
2021-11-17T17:43:59-0500

(a)


"P(X>5000)=e^{-5000/1000}=e^{-5}\approx0.006738"

(b)


"P(X>5000|X>1000)"

"=P(X>4000+1000|X>1000)=P(X>4000)"

"=e^{-4000/1000}=e^{-4}\approx0.018316"